# Learning production functions for supply chains with graph neural networks

### v1

@@ 6 Conclusion
In this work, we have formalized TPGs, with supply chain networks as a canonical example, and developed a new class of GNNs to handle them. Our models successfully achieve two essential objectives in this setting—inferring production functions and predicting future edges—while preexisting GNNs focused on the latter. Furthermore, we build a new simulator, SupplySim, to evaluate models under varied and controlled TPG settings, and release the simulator and synthetic datasets to encourage further research development. In future work, we hope to apply our model to TPGs in other domains, such as biological, and to develop theoretical results for our inventory module, such as establishing identifiability conditions and connecting it to causal inference.
@@ Acknowledgements
S.C. was supported in part by an NSF Graduate Research Fellowship, the Meta PhD Fellowship, and NSF award CCF-1918940. J.L. was supported in part by NSF awards OAC-1835598, CCF-1918940, DMS-2327709, and Stanford Data Applications Initiative. From Hitachi, we are thankful to Arnab Chakrabarti for his continuous support.
Appendix
Appendix A Details about model
We provide additional details about our model architecture and evaluation framework here.
@@ A.1 SC-TGN
Here we describe SC-TGN, our extended version of TGN (Rossi et al.,, 2020), in greater detail. Like TGN, our model consists of the following modules:
@@
Embedding. Finally, the embedding module produces an embedding $\mathbf{z}_{i}^{(t)}$ of node $i$ at time $t$. Following the original work, we explore both an “ID” embedding, where the embedding is simply the node’s memory, and embeddings that are constructed by applying a GNN to the memories. For the GNN, we use the UniMP model (Shi et al.,, 2021), which is the model used in the TGN implementation by Temporal Graph Benchmark (Huang et al.,, 2023). UniMP is a graph transformer architecture with multi-head attention; we refer the reader to the original text for details. The advantage of the GNN-based embedding over the ID embedding is that the GNN allows the node’s embedding to adapt over time, even if the node itself has not been involved in a transaction in a while, thus avoiding the memory staleness problem (Kazemi et al.,, 2020). In Table 5, we show that using a GNN to construct node embeddings from the node memories, instead of directly using the node memories, greatly improves performance of SC-TGN.
@@ A.2 SC-GraphMixer
Here we describe SC-GraphMixer, our extended version of GraphMixer (Cong et al.,, 2023), in greater detail. Like GraphMixer, our model consists of a link encoder and a node encoder:
@@
  • Then, GraphMixer applies a 1-layer MLP-mixer (Tolstikhin et al.,, 2021) to $\mathbf{T}_{i}(t)$:
@@
A.3 Penalties and caps from inventory module
As described in the main text, when our inventory module is combined with a GNN, such as SC-TGN or SC-GraphMixer, we can optionally use the inventory module to help with predicting future edges, in addition to its primary purpose, which is to learn production functions. To help with predicting edge existence, the inventory module penalizes impossible transactions, as described in (8), and adds its penalty to the original prediction $\hat{y}$. To help with predicting edge weight, the inventory module computes a cap, i.e., the maximum amount possible in the transaction, as described in (9), and takes the minimum of its cap and $\hat{y}$. Note that, if the inventory module’s attention weights match the ground-truth production functions and the inventories reflect what the firm truly has at the time of the transaction, these penalties and caps can only help, not hurt, the prediction. This is because the penalty (8) will always be 0 for a true transaction while it can be negative for a false transaction, so adding the penalties can only help. Similarly, the amount cap (9) is guaranteed to be greater than or equal to the true transacted amount, so taking the minimum of the cap and the original predicted amount can also only help.
@@ A.3 Penalties and caps from inventory module
However, even if the learned attention weights match the ground-truth production functions closely, it is not guaranteed that the inventory reflects what each firm truly has at the time of the transaction. One minor case that we see of this is that the inventory is only updated up to time $t$, but the current batch could include transactions at time $t+1$ or later; thus, we only apply penalties or caps to transactions in time $t$, since at time $t+1$ and later, the inventory will look different. A more serious case is when we are systematically missing transactions in the dataset, so as a result, we are underestimating the number of products that each firm has in its inventory, so penalties are too large and caps are too low. We investigate this issue in Appendix C.2 and show that, when some firms are missing from the synthetic data, the penalties and caps indeed hurt edge prediction performance, even if the attention weights are perfectly learned (we set them to the ground-truth production functions). So, it is preferable not to use the penalties and caps in those settings. In future work, it would be useful to explore how to make the inventory module more robust to missing data, so that we do not underestimate the number of products in the firms’ inventories even when transactions are missing. For example, some smoothing could be applied such that when we observe a firm buying product $p$, we also add a bit of $p^{\prime}$ to its inventory, where $p^{\prime}$ is a “similar” product, where similarity could be defined based on co-buying patterns.
@@ A.4 Model training and evaluation
We use mean average precision (MAP) to evaluate how well the inventory module learned the ground-truth production functions. Let $\mathcal{P}_{p_{o}}$ represent the ground-truth set of parts for product $p_{o}$, let $\alpha_{p_{o}}=\{\alpha_{p_{o}p_{1}},\alpha_{p_{o}p_{2}},\cdots,\alpha_{p_{o}p_{m}}\}$ represent the inventory module’s learned attention weights for $p_{o}$, and let $\textrm{pos}(\alpha_{p_{o}},k)$ represent the product at rank $k$ in $\alpha_{p_{o}}$. We define average precision as
$\textrm{AvePrec}(p_{o})$ $=\frac{1}{|\mathcal{P}_{p_{o}}|}\sum_{k=1}^{m}\textrm{Prec@K}(\alpha_{p_{o}},\mathcal{P}_{p_{o}},k)\cdot\mathbbm{1}[\textrm{pos}(\alpha_{p_{o}},k)\in\mathcal{P}_{p_{o}}]$ (17)
@@
In other words, average precision takes the average of Precision@K over all ranks $k$ where there is a relevant item. Average precision and Precision@K are both common metrics in information retrieval, but we prefer average precision here since it can account for varying numbers of relevant items over queries, which applies to our setting since the size of $\mathcal{P}_{p_{o}}$ varies across products.
@@
As described in the main text, we sample two types of negatives, perturbation negatives and historical negatives, to accompany each positive transaction. Temporal Graph Benchmark (Huang et al.,, 2023) also samples both types of negatives, although we modify the procedure slightly for hyperedges. For a positive transaction $e(s,b,p,t)$, we have three types of perturbation negatives: $e(s^{\prime},b,p,t)$, where a random supplier is sampled; $e(s,b^{\prime},p,t)$, where a random buyer is sampled; and $e(s,b,p^{\prime},t)$, where a random product is sampled. A historical negative is a transaction $e(s^{\prime},b^{\prime},p^{\prime},t)$ such that $s^{\prime},b^{\prime},p^{\prime}$ appeared in the train set, but not at time $t$. To make historical negatives even harder, we sample historical negatives proportionally to their count in the train set. For both types of negatives, we ensure that the sampled negative did not actually occur at time $t$. We sample 9 negatives following the perturbation procedure (with 3 of each type) and 9 negatives following the historical procedure. It is possible, though, that more than 9 negatives are historical (if the perturbation procedure results in a historical negative) or more than 9 negatives are perturbations (if a historical negative happens to be a one-node perturbation).
We use mean reciprocal rank (MRR) to evaluate performance for predicting edge existence. For each positive transaction $e$ and its corresponding set of negatives $\mathcal{N}_{e}$ in batch $B$, MRR quantifies the position of the positive transaction among the negative samples:
$MRR=\frac{1}{|B|}\sum_{e\in B}\left(\frac{\sum_{n\in\mathcal{N}_{e}}\mathbbm{1}[\hat{y}_{n}<\hat{y}_{e}]+\sum_{n\in\mathcal{N}_{e}}\mathbbm{1}[\hat{y}_{n}\leq\hat{y}_{e}]}{2}+1\right)^{-1}.$ (19)
Following the Temporal Graph Benchmark (Huang et al.,, 2023), we compute an “optimistic” MRR where we tie-break in favor of the positive transaction and a “pessimistic” MRR where we tie-break in favor of the negative transactions, and take the average of the two. We refer to 0.1 as the random baseline, since we sample 18 negatives in total and, if all transactions received the same score, then the MRR would be $((0+18)/2+1)^{-1}=0.1$. Similarly, if we assigned uniform random scores to all 19 alternatives, the expected rank of the positive is 10 (9 negatives before and 9 negatives after), so the reciprocal is 0.1.
@@
We preprocess transaction amounts by applying log-scaling then standard scaling, where we subtract the mean and divide by the standard deviation, so that amounts now have zero mean and variance of 1. We use the scaled amount as an edge feature, to stabilize training, and as a prediction target, to make RMSE comparable across datasets. However, note that we use the raw amount to update the inventory module, as required in (1)-(2). Then, we compare the predicted amount $\hat{y}$ to the true amount, $y$, using root mean squared error:
@@ Appendix B Supply chains data
Attribute Counts SS-std Tesla IED
@@ B.1 Real-world supply chains data
For our experiments, we constructed two large-scale, real-world supply chain datasets by compiling data from TradeSparq, a third-party data provider. We rely on a wide variety of data sources from authorized government sources across over 60 countries, including authorities like the Automated Manifest System, operated by US Customs and Border Protection, and other government customs departments. The data appear in the form of bills of lading (BoL), receipts of reported transactions, and custom declarations. Each transaction includes information such as data source, buyer title, supplier title, cost of reported transaction, product (in terms of HS code and description), import and export country, amount (weight and quantity unit), among other columns.
We have access to the data through an API, which allows us to make queries with associated costs. Due to these costs, we cannot reconstruct the entire supply chain network, but we focus on certain products and grow the network outwards via a breadth-first-search-like process. The construction begins by selecting product-specific HS codes and identifying prominent Tier-0 supplier firms involved in high-volume trading. The process captures all transactions from Tier-0 suppliers, then aggregates their group companies to identify Tier-1 suppliers, recording all corresponding transactions. This method is repeated to determine Tier-2 suppliers by analyzing the group companies of Tier-1 firms. All transactions between Tier-1 and Tier-2 firms are also collected. This systematic approach ensures a thorough mapping of supplier interactions up to Tier-2, providing detailed insights into the supply chain network. We illustrate our process as a flowchart in Figure 3.
Start Select product HS codes Select high amount trading prominent Tier-0 supplier firms of the product Get all transactions between Tier-0 firms setting as suppliers and their buyers Aggregate group companies of the Tier-0 firms & setting them as buyers get all Tier-1 supplier firms Get all transactions between Tier-0 & Tier-1 firms Aggregate group companies of the Tier-1 firms & setting them as buyers get all Tier-2 supplier firms Get all transactions between Tier-1 & Tier-2 firms End
Figure 3: Overview of supply chain network data construction process.
Automotive dataset.
@@
Limitations of real-world supply chains data.
TradeSparq data provides a rare opportunity to access transaction-level information, while most supply chain studies are done on much more aggregated data (e.g., imports and exports aggregated over country) or, at best, static firm-level networks, without time-varying transactions (Inoue and Todo,, 2019; Carvalho and Tahbaz-Salehi,, 2019; Guan et al.,, 2020). However, there are still limitations of the real-world data, which motivate the need for synthetic data. We list some of these limitations below:
@@
  • Missing production functions. For most firms and products, we are missing their true production functions, which limits our ability to evaluate our inventory module.
  • Proprietary data. Since the TradeSparq data is proprietary, we are not allowed to release it.
@@ B.2 Details about SupplySim
  • We can generate any number of shocks, of varying size and source, and test the model’s ability to predict how transactions change under the shocks. In real data, large shocks are rare and confounded by many other contemporaneous factors so it is difficult to isolate their effect (e.g., shocks during the COVID-19 pandemic).
@@
Thus, SupplySim addresses all the limitations of real-world supply chains data, at the cost of being synthetic, instead of real-world, data. To mitigate this cost, we develop SupplySim to be as realistic as possible: for example, we ensure that the generated networks match real-world networks on key characteristics, such as degree distribution and clustering, and we use a realistic agent-based model from economics to generate time-varying transactions over the network.
@@ B.2.1 Static graphs
Based on the assignment of suppliers to products, and the existing production graph $G_{\textrm{prod}}$, we can now determine which input products each firm needs to buy in order to make the products it supplies. We iterate through pairs $(b,p)$, where firm $b$ needs to buy product $p$, and select a supplier firm $s$ from all suppliers of $p$, with probability proportional to $s$’s current number of buyers. This assignment mechanism, known as preferential attachment (Newman,, 2001), reflects a rich-get-richer dynamic that results in power law degree distributions, which are known to appear in real-world supply chain networks (Fujiwara and Aoyama,, 2010; Inoue and Todo,, 2019; Zhao et al.,, 2019). Real-world firm-firm networks are also reported to have disassortativity (negative correlation between the degrees of connected nodes) and low clustering, which we also observe in our generated firm-firm networks: the degree assortativity is -0.253 and average clustering coefficient is 0.246, which is similar to the average clustering coefficient of 0.179 reported in Zhao et al., (2019).
B.2.2 Generating time-varying transactions
Now, given the static graphs, we want to generate time-varying transactions between firms. We base our simulation model on the ARIO model, an agent-based model widely used in economics to simulate propagation over supply chains (Hallegatte,, 2008; Inoue and Todo,, 2019; Guan et al.,, 2020). At any point in the simulation, we keep track of each firm’s inventory $\mathbf{x}_{i}^{(t)}$; all incomplete orders $\mathcal{I}^{(t)}$, where an order is defined as $o(b,s,p,k)$, from buyer firm $b$ to supplier firm $s$ for $k$ amount of product $p$; and all orders that were newly completed at time $t$, $\mathcal{C}^{(t)}$. Then, at each timestep $t$, we iterate through firms, with each firm $f$ taking the following actions.
@@ B.2.2 Generating time-varying transactions
Figure 5: Visualizing the number of transactions per timestep, over the three synthetic datasets: SS-std, SS-shocks, and SS-missing.
@@
In our synthetic data with shocks, SS-shocks, we use $p_{\textrm{shock}}=0.01$, resulting in 10 shocks over 200 timesteps: shocks to product 0 at times 32 and 144, product 2 at times 42 and 155, product 3 at times 51, 90, and 144, and product 4 at times 16, 58, and 131. We allow a recovery rate of $r=1.25$, resulting in a recovery time of approximately 30 timesteps from the initial shock to full recovery (since we set the stable supply to 1000x of $supply_{\textrm{shock}}$). In Figure 5, we visualize the number of transactions over time in each synthetic data (SS-std, SS-shocks, and SS-missing). While the missing firms reduce the number of transactions by a consistent fraction, the shocks dramatically disrupt the pattern of transactions, slowing them down whenever there is a shortage then trying to catch up whenever the Tier 0 products have recovered.
@@
Appendix C Details about experiments
Here we provide details about our experimental set-up, including our baselines and model training / testing procedures.

### v3

@@ 6 Discussion
In this work, we have formalized temporal production graphs (TPGs), with supply chains as a canonical example, and developed a new class of GNNs that can handle TPGs. We also release a new open-source simulator, SupplySim, which enables rigorous model testing and future research development in this area. Our models successfully achieve two essential objectives—inferring production functions and predicting future edges—while preexisting GNNs focused on the latter. Our models can be used in a wide variety of real-world scenarios, such as demand forecasting, early risk detection, and inventory optimization. To concretely illustrate these possibilities, we conducted a case study, where we showed how the model can be used to generate future transactions under supply chain shocks and disruptions. In future work, we hope to explore other potential use cases in supply chains, apply our model to TPGs in other domains, and develop theoretical results for our inventory module, such as establishing identifiability conditions and connecting it to causal inference.
@@ Acknowledgements
S.C. was supported in part by an NSF Graduate Research Fellowship, the Meta PhD Fellowship, and NSF award CCF-1918940. J.L. was supported in part by NSF awards OAC-1835598, CCF-1918940, DMS-2327709, and Stanford Data Applications Initiative. The authors thank Emma Pierson and members of Jure Leskovec’s lab for helpful comments on the draft. From Hitachi, we are thankful to Arnab Chakrabarti for his continuous support.
Appendix A Model Architecture and Evaluation
In this section, we provide additional details about our SC-TGN architecture (A.1), our SC-GraphMixer architecture (A.2), our inventory module’s link prediction penalties (A.3), and our evaluation framework (A.4).
@@ A.1 SC-TGN
Here we describe SC-TGN, our extended version of TGN (Rossi et al. 2020), in greater detail. Like TGN, our model consists of the following modules:
@@
Embedding. Finally, the embedding module produces an embedding $\mathbf{z}_{i}^{(t)}$ of node $i$ at time $t$. Following the original work, we explore both an “ID” embedding, where the embedding is simply the node’s memory, and embeddings that are constructed by applying a GNN to the memories. For the GNN, we use the UniMP model (Shi et al. 2021), which is the model used in the TGN implementation by Temporal Graph Benchmark (Huang et al. 2023). UniMP is a graph transformer architecture with multi-head attention; we refer the reader to the original text for details. The advantage of the GNN-based embedding over the ID embedding is that the GNN allows the node’s embedding to adapt over time, even if the node itself has not been involved in a transaction in a while, thus avoiding the memory staleness problem (Kazemi et al. 2020). In Table 7, we show that using a GNN to construct node embeddings from the node memories, instead of directly using the node memories, greatly improves performance of SC-TGN.
@@ A.2 SC-GraphMixer
Here we describe SC-GraphMixer, our extended version of GraphMixer (Cong et al. 2023), in greater detail. Like GraphMixer, our model consists of a link encoder and a node encoder:
@@
  • Then, GraphMixer applies a 1-layer MLP-mixer (Tolstikhin et al. 2021) to $\mathbf{T}_{i}(t)$:
@@
A.3 Penalties from inventory module
As described in the main text, when our inventory module is combined with a GNN, such as SC-TGN or SC-GraphMixer, we can use the inventory module to not only learn production functions but also help with predicting future edges. To help with predicting edge existence, the inventory module penalizes impossible transactions, i.e., some $(s,b,p,t)$ where the supplier $s$ does not have enough parts to make product $p$. We define the penalty as
Then, our model’s new prediction for edge existence becomes the sum of the original $\hat{y}$ in (7) and the penalty (14). To help with predicting edge weight, i.e., transaction amount, the inventory module computes a cap on the maximum amount of product $p$ that supplier $s$ could produce at this time.
$cap(s,b,p,t)$ $=\min_{p^{\prime}\in[m];\alpha_{pp^{\prime}}>0}\{\frac{\mathbf{x}_{s}^{(t)}[p^{\prime}]}{\alpha_{pp^{\prime}}}\}$ (15)
Then, our new predicted amount is the minimum of the original predicted amount (7) and the cap (after applying the same log-scaling and standard scaling to the cap that we did to all amounts). Note that, if $\alpha_{pp^{\prime}}$ matches the ground-truth production function $\mathcal{F}_{p}$, i.e., exactly $\alpha_{pp^{\prime}}$ amount of $p^{\prime}$ is needed to make one unit of $p$, and the inventory at time $t$ is correct, these penalties and caps can only help, not hurt, the prediction. Specifically, the penalty will always be 0 for a true transaction while it can be negative for a false transaction, and the cap is guaranteed to be greater than or equal to the true transacted amount. Since the inventory is only updated up to time $t$, but the current batch could include transactions at time $t+1$ or later, we only apply penalties or caps to transactions in time $t$, since at time $t+1$ and later, the inventory will look different.
@@ A.3 Penalties from inventory module
However, even if the learned attention weights match the ground-truth production functions closely, it is not guaranteed that the inventory reflects what each firm truly has at the time of the transaction. One minor case that we see of this is that the inventory is only updated up to time $t$, but the current batch could include transactions at time $t+1$ or later; thus, we only apply penalties and caps to transactions in time $t$, since at time $t+1$ and later, the inventory will look different. A more serious case is when we are systematically missing transactions in the dataset, so as a result, we are underestimating the number of products that each firm has in its inventory, so penalties are too large and caps are too low. We investigate this issue in Appendix C.2 and show that, when some firms are missing from the synthetic data, the penalties and caps indeed hurt edge prediction performance, even if the attention weights are perfectly learned (we set them to the ground-truth production functions). So, it is preferable not to use the penalties and caps in those settings. In future work, it would be useful to explore how to make the inventory module more robust to missing data, so that we do not underestimate the number of products in the firms’ inventories even when transactions are missing. For example, some smoothing could be applied such that when we observe a firm buying product $p$, we also add a bit of $p^{\prime}$ to its inventory, where $p^{\prime}$ is a “similar” product, where similarity could be defined based on co-buying patterns.
@@ A.4 Model training and evaluation
We use mean average precision (MAP) to evaluate how well the inventory module learned the ground-truth production functions. Let $\mathcal{P}_{p}$ represent the ground-truth set of parts for product $p$, let $\alpha_{p}=\{\alpha_{pp_{1}},\alpha_{pp_{2}},\cdots,\alpha_{pp_{m}}\}$ represent the inventory module’s learned attention weights for $p$, and let $\alpha_{p}^{(k)}$ represent the product at rank $k$ in $\alpha_{p}$. We define average precision as
$\textrm{AvePr}(p)$ $=\frac{1}{|\mathcal{P}_{p}|}\sum_{k=1}^{m}\textrm{Pr@K}(\alpha_{p},\mathcal{P}_{p},k)\cdot\mathds{1}[\alpha_{p}^{(k)}\in\mathcal{P}_{p_{o}}]$ (17)
@@
In other words, average precision takes the average of Precision@K over all ranks $k$ where there is a relevant item. Average precision and Precision@K are both common metrics in information retrieval, but we prefer average precision here since it can account for varying numbers of relevant items over queries, which applies to our setting since the size of $\mathcal{P}_{p}$ varies across products.
@@
As described in the main text, we sample two types of negatives, perturbation negatives and historical negatives, to accompany each positive transaction. Temporal Graph Benchmark (Huang et al. 2023) also samples both types of negatives, although we modify the procedure slightly for hyperedges. For a positive transaction $e(s,b,p,t)$, we have three types of perturbation negatives: $e(s^{\prime},b,p,t)$, where a random supplier is sampled; $e(s,b^{\prime},p,t)$, where a random buyer is sampled; and $e(s,b,p^{\prime},t)$, where a random product is sampled. A historical negative is a transaction $e(s^{\prime},b^{\prime},p^{\prime},t)$ such that $s^{\prime},b^{\prime},p^{\prime}$ appeared in the train set, but not at time $t$. Compared to a uniform random negative, these negatives are far harder to distinguish from the positive transaction. For the perturbations, they match the positive transaction on two of the three nodes, requiring the model to make a distinction based on the third node. The historical negatives are even harder: this transaction has been observed before, but it does not occur at this timestep, requiring the model to learn time-varying properties of the graph. To make historical negatives even harder, we sample historical negatives proportionally to their count in the train set. For both types of negatives, we ensure that the sampled negative did not actually occur at time $t$.
We sample 9 negatives following the perturbation procedure (with 3 of each type) and 9 negatives following the historical procedure. It is possible, though, for more than 9 negatives to be historical (if the perturbation procedure results in a historical negative) or more than 9 negatives to be perturbations (if a historical negative happens to be a one-node perturbation). We use mean reciprocal rank (MRR) to evaluate performance for predicting edge existence. For each positive transaction $e$ and its corresponding set of negatives $\mathcal{N}_{e}$, MRR computes the rank of the positive transaction’s predicted score, $\hat{y}_{e}$, among the negative samples’ predicted scores. Following the Temporal Graph Benchmark (Huang et al. 2023), we compute an “optimistic” rank, $r_{\textrm{opt}}(e)$, where we tie-break in favor of the positive transaction, and a “pessimistic” rank, $r_{\textrm{pes}}(e)$, where we tie-break in favor of the negative transactions:
$r_{\textrm{opt}}(e)$ $=\sum_{n\in\mathcal{N}_{e}}\mathds{1}[\hat{y}_{n}<\hat{y}_{e}]$ (19)
$r_{\textrm{pes}}(e)$ $=\sum_{n\in\mathcal{N}_{e}}\mathds{1}[\hat{y}_{n}\leq\hat{y}_{e}].$ (20)
Then, MRR takes the reciprocal of the average of the two ranks (+1 to avoid dividing by 0), over all positive transactions $e$ in batch $B$:
$\textrm{MRR}=\frac{1}{|B|}\sum_{e\in B}\left(\frac{r_{\textrm{opt}}(e)+r_{\textrm{pos}}(e)}{2}+1\right)^{-1}.$ (21)
We refer to 0.1 as the random baseline, since we sample 18 negatives in total and, if all transactions received the same score, then the MRR would be $((0+18)/2+1)^{-1}=0.1$. Similarly, if we assigned uniform random scores to all 19 alternatives, the expected rank of the positive is 10 (9 negatives before and 9 negatives after), so the reciprocal is 0.1.
@@
We preprocess transaction amounts by applying log-scaling then standard scaling, where we subtract the mean and divide by the standard deviation, so that amounts now have zero mean and variance of 1. Scaling is useful when using amount as an edge feature, to stabilize training, and as a prediction target, to make RMSE comparable across datasets. However, note that we use the raw amount to update the inventory module, as required in (1)-(2). Then, we compare the predicted amount $\hat{y}$ to the true amount, $y$, using root mean squared error:
@@ Appendix B Supply Chains Data
In this section, we provide greater detail about our real-world supply chains datasets from TradeSparq (B.1), our simulator SupplySim for generating realistic supply chains data (B.2), and network characteristics of our synthetic data, especially how they relate to real-world supply chain networks (B.3). We also provide a summary of dataset statistics in Table 4.
SS-std Tesla IED
@@ B.1 Real-world supply chains data
For our experiments, we constructed two large-scale, real-world supply chain datasets by compiling data from TradeSparq, a third-party data provider. We rely on a wide variety of data sources from authorized government sources across over 60 countries, including authorities like the Automated Manifest System, operated by US Customs and Border Protection, and other government customs departments. The data appear in the form of bills of lading, receipts of reported transactions, and custom declarations. Each product being transacted is described with a Harmonized System (HS) code, where HS6 is an internationally recognized six digit number that describes products in terms of their chapters (first two digits), headings (next two digits), and subheadings (last two digits). For example, the HS code 840731 refers to the chapter 84 (chapter: Machinery), 07 (heading: Spark-ignition Reciprocating Piston Engines), and 31 (sub-heading: engines type not over 50CC cylinder capacity). Beside HS code, each transaction also includes data source, the buyer firm’s name and ID, the supplier firm’s name and ID, a product description in free text, import and export country, cost of reported transaction, weight, and number of units, among other columns.
We have access to the data through an API, which allows us to make queries with associated costs. Due to these costs, we cannot reconstruct the entire supply chain network, but we focus on certain products and grow the network outwards via a breadth-first-search-like process. The construction begins by selecting product-specific HS codes and identifying prominent Tier-0 supplier firms involved in high-volume trading. The process captures all transactions from Tier-0 suppliers, then aggregates their group companies to identify Tier-1 suppliers, recording all corresponding transactions. This method is repeated to determine Tier-2 suppliers by analyzing the group companies of Tier-1 firms. All transactions between Tier-1 and Tier-2 firms are also collected. This systematic approach ensures a thorough mapping of supplier interactions up to Tier-2, providing detailed insights into the supply chain network.
Tesla dataset.
@@
Limitations of real-world data.
TradeSparq data provides a rare opportunity to access transaction-level information, while most supply chain studies are done on much more aggregated data (e.g., imports and exports aggregated over country) or, at best, static firm-level networks, without time-varying transactions (Inoue and Todo 2019; Carvalho and Tahbaz-Salehi 2019; Guan et al. 2020). However, a key limitation of the real-world data is that we cannot release it, which complicates replicability and opportunities for future work. Furthermore, the data itself has limitations that arise from the complexity of real-world supply chains:
@@
  • Missing production functions. For most firms and products, we are missing their true production functions, which we need to evaluate our inventory module.
@@ B.2 Details about SupplySim
  • We can inject any number of shocks, of varying size and source, and test the model’s ability to predict how transactions change under the shocks. In real data, large shocks are rare and confounded by many other contemporaneous factors so it is difficult to isolate their effect (e.g., shocks during the COVID-19 pandemic).
@@
Thus, SupplySim addresses all the limitations of real-world supply chains data, at the cost of being synthetic, instead of real-world, data. To mitigate this cost, we develop SupplySim to be as realistic as possible: for example, we ensure that the generated networks match real-world networks on key characteristics, such as degree distribution and clustering, and we use a realistic agent-based model to generate time-varying transactions over the network.
@@
Based on the assignment of suppliers to products, and the existing production graph $G_{\textrm{prod}}$, we can now determine which input products each firm needs to buy in order to make the products it supplies. We iterate through pairs $(b,p)$, where firm $b$ needs to buy product $p$, and select a supplier firm $s$ from all suppliers of $p$, with probability proportional to $s$’s current number of buyers. This assignment mechanism, known as preferential attachment (Newman 2001), reflects a rich-get-richer dynamic that results in power law degree distributions, which are known to appear in real-world supply chain networks (Fujiwara and Aoyama 2010; Inoue and Todo 2019; Zhao, Zuo, and Blackhurst 2019).
Generating time-varying transactions.
Now, given the static graphs, we want to generate time-varying transactions between firms. We base our simulation model on the ARIO model, an agent-based model widely used in economics to simulate propagation over supply chains (Hallegatte 2008; Inoue and Todo 2019; Guan et al. 2020). At any point in the simulation, we keep track of each firm’s inventory $\mathbf{x}_{i}^{(t)}$; all incomplete orders $\mathcal{I}^{(t)}$, where an order is defined as $o(b,s,p,k)$, from buyer firm $b$ to supplier firm $s$ for $k$ amount of product $p$; and all orders that were newly completed at time $t$, $\mathcal{C}^{(t)}$. Then, at each timestep $t$, we iterate through firms, with each firm $f$ taking the following actions.
@@
In our synthetic data with shocks, SS-shocks, we use $p_{\textrm{shock}}=0.01$, resulting in 10 shocks over 200 timesteps: shocks to product 0 at times 32 and 144, product 2 at times 42 and 155, product 3 at times 51, 90, and 144, and product 4 at times 16, 58, and 131. We allow a recovery rate of $r=1.25$, resulting in a recovery time of approximately 30 timesteps from the initial shock to full recovery (since we set the stable supply to 1000x of $supply_{\textrm{shock}}$). In Figure 2d, we visualize the number of transactions over time in each synthetic data (SS-std, SS-shocks, and SS-missing). While the missing firms reduce the number of transactions by a consistent fraction, the shocks dramatically disrupt the pattern of transactions, slowing them down whenever there is a shortage then trying to catch up whenever the Tier 0 products have recovered.
@@
B.3 Characteristics of real and synthetic data
In Figure 2, we visualize several characteristics of the synthetic data generated by SupplySim, which match known characteristics of real-world supply chain networks.
Degree distribution.
Supply chain networks are known to exhibit power law degree distributions in firm-firm networks: a few large firms dominate supply for many products (Fujiwara and Aoyama 2010; Inoue and Todo 2019; Zhao, Zuo, and Blackhurst 2019). In Figure 2a, we show the degree distribution from our synthetic data, after we construct a firm-firm network, where an edge from firm $i$ to $j$ indicates that firm $i$ is a supplier of firm $j$. As expected, the degrees in our synthetic data also follow a power law distribution due to our preferential attachment mechanism for assigning buyers to suppliers (i.e., probability proportional to how many buyers they already have).
Community structure.
Supply chains are also known to have community structure, corresponding to industrial sectors or geographical regions (Fujiwara and Aoyama 2010). Following Fujiwara and Aoyama (2010), we quantify community structure with modularity, denoted as $Q$, which assesses the quality of a network partition into communities:
$Q=\frac{1}{2E}\sum_{ij}(A_{ij}-\frac{k_{i}k_{j}}{2E})\mathds{1}[c_{i}=c_{j}],$ (28)
where $E$ is the number of edges in the network; $A_{ij}$, as the adjacency matrix, indicates whether nodes $i$ and $j$ are connected; $k_{i}$ is the degree of node $i$; and $\mathds{1}[c_{i}=c_{j}]$ indicates whether $i$ and $j$ are in the same community within the partition. Intuitively, modularity compares the number of observed edges $(A_{ij})$ to the number of expected edges ($k_{i}k_{j}/2E$) between nodes in the same community. Modularity can range from -0.5 to 1, and any positive value indicates community structure, meaning that there are more observed edges than expected within communities.
We try both the Louvain algorithm (Blondel et al. 2008) and a greedy algorithm (Clauset, Newman, and Moore 2004), the same used by Fujiwara and Aoyama (2010), to partition our firm-firm network into communities. Under both algorithms, we find that modularity is strongly positive, indicating community structure in our synthetic networks: $Q=0.35$ with the Louvain algorithm and $Q=0.37$ with the greedy algorithm. In Figure 2b, we visualize the five communities found by the Louvain algorithm, and we color each firm by the average tier of its products supplied, from Tier 0 (raw materials) to Tier 5 (consumer products). From these colors, we can see clear patterns in the communities, as the firms within them tend to specialize in products of similar tiers. This structure arises naturally from how we constructed our synthetic data, since firms were restricted to supplying products within two consecutive tiers.
Clustering.
While supply chains have community structure, they also have low clustering, since unlike social networks, a firm’s supplier’s supplier is unlikely to also be this firm’s supplier (Fujiwara and Aoyama 2010; Zhao, Zuo, and Blackhurst 2019). In other words, we should expect fewer triangles in this network. A node’s clustering coefficient is defined as the number of observed edges between its neighbors, divided by the total possible number of edges between its neighbors, which is $k_{i}(k_{i}-1)/2$. In Figure 2c, we show the distribution of clustering coefficients over nodes in the firm-firm network. Most nodes have low clustering, with an average clustering coefficient of 0.246.
Disassortativity.
Another unique characteristic of supply chain networks is that they exhibit disassortativity, meaning high-degree nodes tend to be connected to low-degree nodes, instead of to other high-degree nodes (Fujiwara and Aoyama 2010; Inoue and Todo 2019). Formally, assortativity is measured as the Pearson correlation between the degrees of connected nodes (Newman 2002), where a negative correlation is interpreted as disassortativity. We measure assortativity on our firm-firm networks and find a correlation of $r=-0.253$, indicating disassortativity.
Appendix C Experimental Details
Here we provide details about our experimental set-up, including our baselines and model training / testing procedures. Our experiments were run on a machine with 768 GB of memory and 10 GPUs (NVIDIA GeForce RTX 2080 Ti). Its operating system is Ubuntu 16.04. Our code to run all experiments is available at https://github.com/snap-stanford/supply-chains/tree/tgb, including necessary software libraries and versions.